OpenStudy (brucebaner):

What is the value of the expression? A. 9^−48 B. 9^−2 C. 9^2 D. 9^22

OpenStudy (brucebaner):

@DanJS

OpenStudy (brucebaner):

i think its C

OpenStudy (danjs):

Here you can apply just those couple rules we discussed already....

OpenStudy (danjs):

\[\frac{ 9^8 }{ 9^{-4}*9^{10} } = \frac{ 9^8 }{ 9^{-4+10} } = \frac{ 9^8 }{ 9^6 }\]

OpenStudy (danjs):

i just combined the powers on the like base in the bottom so far...

OpenStudy (danjs):

then recall, to move the bottom term to the top, you have to switch the sign on the power to negative

OpenStudy (danjs):

the final answer will be 9^(8-6)

OpenStudy (danjs):

so you were right!